### v1

@@ Abstract.
In this work, we develop the method of multipliers for electromagnetic Dirac operators and establish sufficient conditions on the magnetic and electric fields that guarantee the absence of point spectrum. Our results can be viewed as a generalization of those obtained by Cossetti, Fanelli and Krejčiřík (2020), which treated the purely magnetic case. We also emphasize that, in the massless case, our approach covers Coulomb-type potentials of the form $V(x)=\frac{1}{|x|}\big{(}\nu\mathbb{I}+\mu\beta+i\delta\beta\big{(}\boldsymbol{\alpha}\cdot\frac{x}{|x|}\big{)}\big{)}.$
@@ 1. Introduction
Here $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3}),$ with $\alpha_{j},$ $j=1,2,3$ and $\beta$ being the standard $4\times 4$ Hermitian Dirac matrices, $m\geq 0$ denotes the mass of the particle, $V$ is a potential function that associates to each $x\in\mathbb{R}^{3}$ a $4\times 4$ Hermitian matrix $V(x).$ Moreover $\nabla_{\!\text{A}}$ stands for the magnetic gradient $\nabla-iA,$ where $A\colon\mathbb{R}^{3}\to\mathbb{R}^{3}$ is a magnetic vector potential. The operator (1.1) represents the Hamiltonian of a relativistic particle of spin-$\tfrac{1}{2}$ under the action of an electric field $V$ and a magnetic field $B=\nabla\times A.$
Compared to the cases of purely electric [23, 28, 25, 10, 15, 32, 33] or purely magnetic [2, 19, 18, 6, 22] perturbations, the literature available related to the absence of point spectrum of electromagnetic Dirac operators is much scarcer and much less complete. The results presented in this work are motivated by the need to fill in this gap. One of the first works available on this topic is due to Roze [31], who considered electric potentials of the form $V(x)=v(x)\mathbb{I},$ and proved that if $A$ and $v$ are smooth and satisfy
@@
then the Dirac operator has no eigenvalues embedded in the continuous spectrum. A similar result was proved by Kalf in [24] where he replaced the decay assumption (1.2) by an analogous assumption involving the magnetic field $B$ instead of the magnetic potential $A$. Thus giving a gauge invariant condition for the absence of eigenvalues. The result of Kalf was later extended to more general matrix-valued potentials in [3]. In a more recent result [30] the authors investigated on spectral properties of Dirac operators with magnetic fields of constant direction and Coulomb-type electric potentials.
@@
In the present paper, we will develop the so-called method of multipliers in a general setting in order to establish a source of sufficient conditions (decay/Hardy-type) on the potentials to guarantee the absence of point spectrum (including possibly embedded eigenvalues) of the Dirac Hamiltonian $H_{m}(A,V)$ in dimension $d\geq 3.$ See [8] for a recent survey on the method of multipliers in spectral theory.
@@
In the last decade, starting from the pioneering work [17], this method was shown to be a very powerful tool in spectral theory and, in particular, in proving absence of eigenvalues (embedded and discrete) of very different Hamiltonians [6, 9, 8, 7, 16, 17, 1, 22]. Despite its robustness, this method does not seem to apply directly to general Dirac operators. Indeed, some technical difficulties arise due to the lack of positivity of certain commutators, which is also related to the lack of lower semi-boundedness of the operator itself. See references  [11, 12, 14, 15, 21, 27] for related results obtained using alternative techniques. To recover such positivity, a natural strategy is to employ the well-known supersymmetric structure of the free Dirac operator, $H_{m}(0,0)=-i\boldsymbol{\alpha}\cdot\nabla+m\beta$. This means that squaring out the operator gives
$H_{m}(0,0)^{2}=\begin{pmatrix}H_{P}+m^{2}\mathbb{I}_{\mathbb{C}^{2}}&0\\ 0&H_{P}+m^{2}\mathbb{I}_{\mathbb{C}^{2}},\end{pmatrix}$
@@
where $H_{P}:=-\Delta\mathbb{I}_{\mathbb{C}^{2}}+\boldsymbol{\sigma}\cdot B$ denotes the Pauli operator in $L^{2}(\mathbb{R}^{3})^{2}.$ Here $\sigma=(\sigma_{1},\sigma_{2},\sigma_{3})$ are the standard $2\times 2$ Hermitian Pauli matrices. An important consequence of this spectral relation is that proving absence of eigenvalues of $H_{m}(0,0)$ reduces to prove absence of eigenvalues of second order operators of Schrödinger type for which the aforementioned positivity is guaranteed. It is easy to see that the supersymmetric property is maintained for a purely magnetic Dirac operator, namely for $H_{m}(A,0)$ (see [8] and [22]). However, introducing a general electric potential $V$ breaks this property, making the analysis of such operators much harder. Overcoming this difficulty is precisely the aim of the present work.
@@
The first result of this paper shows that, if the potential satisfies a suitable Hardy-type smallness condition, no eigenvalues are created in the spectral gap.
@@
hold true for every $\psi\in H^{1}(\mathbb{R}^{d})$. Then $\sigma_{p}(H_{m}(A,V))=\varnothing$.
@@
Let $d\geq 3.$ Suppose that $V:\mathbb{R}^{d}\to\mathbb{C}^{N\times N}$ is such that $V\in C^{\infty}(\mathbb{R}^{d}\setminus\{x_{0}\})$ and that $V$ is bounded outside any ball containing the origin. If $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4}\geq 0$ satisfy (1.4) and $B$ and $V$ satisfy (1.5)-(1.8), then $\sigma_{p}(H_{m}(A,V))=\varnothing.$
@@
Let $d\geq 3.$ Suppose that $V$ is as in Theorem 1.2. If $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4}$ satisfy (1.4) and $B$ and $V$ satisfy (1.5)-(1.8), then for any unitary matrix $P\in\mathbb{C}^{N\times N}$ one has $\sigma_{p}(H_{m}(A,PVP^{\ast}))=\varnothing,$
@@
This provides a sufficient condition for assumptions (1.5), (1.6) and (1.8) to be satisfied, i.e. $|W|\approx\mu|x|^{-2},$ for some suitably small coupling constant $\mu>0.$ Notice that in the specific cases, namely (1.5), (1.6) and (1.8), one has $|W|=|x|^{2}|B|^{2},$ $|W|=|x|^{2}|\nabla V|^{2},$ and $|W|=|x||\nabla V|,$ respectively.
On the other hand, one can prove (see [20, Thm. 2.1] and [35]) that if $W$ satisfies a Hardy-type inequality (1.10), then there exist $c,R>0$ such that
@@
Finally observe that, as a consequence of the diamagnetic inequality (see, for instance, [26])
@@
here $\langle x\rangle:=(1+|x|^{2})^{1/2}$ is the standard notation for Japanese brackets. The alternative condition (1.11) shows that in the massive case, or if ${\beta,V}\neq 0$, we can indeed allow for Coulomb-type decay near the origin, but not at infinity.
@@
is included, where the potentials $\mathbb{V}_{\text{el}},\mathbb{V}_{\text{sc}},\mathbb{V}_{\text{am}}$ are customarily called electric, scalar and anomalous magnetic, respectively and which are known to be perturbations that leave invariant the partial wave subspaces.
We emphasize that the impossibility to cover Coulomb-type potentials (1.12) in the massive case is not just technical, indeed it is known that, as soon as $m\neq 0,$ there are Coulomb-type perturbations that do produce discrete eigenvalues in the spectral gap (m,m)(-m,m) (see [13] and references therein for the case $A=0$ and $\mathbb{V}_{\text{el}}$ and [34] for the case $\mathbb{V}_{\text{sc}}$). The situation is different for anomalous magnetic potentials $\mathbb{V}_{\text{am}}.$ Indeed, in this case, since $\{\beta,\mathbb{V}_{\text{am}}\}=0,$ assumption (1.8) is no longer needed, thus Theorem 1.2 is valid for $V$ as in (1.12) with $\mathbb{V}=\mathbb{V}_{\text{am}}$ also in the massive case, namely for any $m\geq 0.$ We emphasize that this does not contradict the explicit eigenvalues and eigenfunctions obtained in [4] for Dirac operators with purely anomalous magnetic perturbations. Indeed, as can be seen from the proof of their result, if the coupling constant $\delta$ (there called $\lambda$) is sufficiently small, then the corresponding eigenfunctions found there are not $L^{2}$-integrable.
@@
This is a generalization of the special case (1.12) with (1.13), and has been extensively studied as it is an interesting model in quantum mechanics (see, for example [4] for further information).
@@
and assume (1.5), (1.6) and, if $m\neq 0,$ assume also (1.8). Moreover assume that
@@
    We refer to [5] for a complete survey on the concept of magnetic field in any dimensions and its definition in terms of differential forms and tensor fields.
@@ 2. Definition of Dirac operator
In this section we gather some basic properties on the Dirac operators. We refer to [36] for a more exhaustive dissertation.
@@
for $j,k\in\{1,2,\dots,d\}$. For an explicit construction of these matrices in any dimension see for example [6]. It is well known that this operator is self-adjoint in $L^{2}(\mathbb{R}^{d})$ with domain $D(H_{m})=H^{1}(\mathbb{R}^{d})$ and has pure continuous spectrum, namely
@@
for $a>0$ small enough (see [36, Sec.4.3]). For the electric potential we shall assume that $V(x)$ is Hermitian for almost all $x\in\mathbb{R}^{d}$. Furthermore, to ensure the self-adjointness of $H_{m}(A,V)$ and, in consequence, that its spectrum is real, we will assume that $V$ is $H_{m}(A)$-bounded with relative bound less than 1. This means that there exist constants $a,b>0$ with $a\leq 1$ such that for every $\psi\in H^{1}(\mathbb{R}^{d})$ the following inequality
@@
holds. This allows us to define $H_{m}(A,V)$ as a self-adjoint operator in $L^{2}(\mathbb{R}^{d})$ with domain $H^{1}(\mathbb{R}^{d})$ by the Kato-Rellich theorem [29, Sec.X.2]. Notice that in some specific cases of special interest, condition (2.2) can be replaced by weaker assumptions.

### v2

@@ Abstract.
In this work, we develop the method of multipliers for electromagnetic Dirac operators and establish sufficient conditions on the magnetic and electric fields that guarantee the absence of point spectrum. In the massless case, our approach covers Coulomb-type potentials of the form $V(x)=\frac{1}{|x|}\big(\nu\mathbb{I}+\mu\beta+i\delta\beta\big(\boldsymbol{\alpha}\cdot\frac{x}{|x|}\big)\big).$ We also adapt the method to show absence of embedded eigenvalues above a threshold which depends on the asymptotic behaviour of the magnetic and electric fields.
@@ 1. Introduction
Here $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3}),$ with $\alpha_{j},$ $j=1,2,3$ and $\beta$ being the standard $4\times 4$ Hermitian Dirac matrices, $m\geq 0$ denotes the mass of the particle, and $V$ is a potential function that associates to each $x\in\mathbb{R}^{3}$ a $4\times 4$ Hermitian matrix $V(x).$ Moreover, $\nabla_{\!\text{A}}$ stands for the magnetic gradient $\nabla-iA,$ where $A\colon\mathbb{R}^{3}\to\mathbb{R}^{3}$ is a magnetic vector potential. The operator (1.1) represents the Hamiltonian of a relativistic particle of spin-$\tfrac{1}{2}$ under the action of an electric field $V$ and a magnetic field $B=\nabla\times A.$
Compared to the cases of purely electric [Cuenin17, Fanelli2019, Kalf1976, Kalf2003, Okaji2003, Schmidt2010, Schmidt2015] or purely magnetic [Balinsky2001, Cossetti2020, Frank2022a, Frank2022, Hundertmark2024] perturbations, the literature available related to the absence of point spectrum of electromagnetic Dirac operators is much scarcer and much less complete. The results presented in this work are motivated by the need to fill in this gap. One of the first works available on this topic is due to Roze [Roze1970], who considered electric potentials of the form $V(x)=v(x)\mathbb{I},$ and proved that if $A$ and $v$ are smooth and satisfy
@@
then the Dirac operator has no eigenvalues embedded in the continuous spectrum. A similar result was proved by Kalf in [Kalf1981], where he replaced the decay assumption (1.2) by an analogous assumption involving the magnetic field $B$ instead of the magnetic potential $A$, thus giving a gauge invariant condition for the absence of eigenvalues. The result of Kalf was later extended to more general matrix-valued potentials in [Berthier1987]. In a more recent result [Richard2007], the authors investigated on spectral properties of Dirac operators with magnetic fields of constant direction and Coulomb-type electric potentials.
More recently, Cossetti, Fanelli and Krejčiřík [Cossetti2020] have proved the total absence of eigenvalues for purely magnetic Dirac operators in dimension $d\geq 3$ if the following inequality
$\int|x|^{2}|B|^{2}|\psi|\leq c\int|\nabla_{\!\text{A}}\psi|^{2},$
holds for every $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$ and for $c>0$ small enough, that is, if the magnetic field is small in a Hardy sense. In addition, Hundertmark and Kovařík [Hundertmark2024], using similar methods, proved the absence of embedded eigenvalues under some weaker hypothesis (See Remark 1.8). These works rely heavily on the supersymmetric structure of the purely magnetic Dirac operator, a property that no longer holds for more general electromagnetic perturbations.
@@
In the present paper, in order to deal with such general perturbations, we will develop the so-called method of multipliers and establish a source of sufficient conditions (decay/Hardy-type) on the potentials to guarantee the absence of point spectrum (including possibly embedded eigenvalues) of the Dirac Hamiltonian $H_{m}(A,V)$ in dimension $d\geq 3.$
@@
In the last decade, starting from the pioneering work [Fanelli2018a], this method was shown to be a very powerful tool in spectral theory and, in particular, in proving absence of eigenvalues (embedded and discrete) of very different Hamiltonians [Hundertmark2020, Cossetti2020, Cossetti2024a, Cossetti2024, Cossetti2020a, Fanelli2018, Fanelli2018a, Hundertmark2024] (see also [Cossetti2024] for a survey on the method of multipliers in spectral theory). Despite its robustness, this method does not seem to apply directly to general Dirac operators. Indeed, several technical difficulties arise from the lack of positivity of certain commutators, a feature closely connected to the unboundedness from below of the operator itself and to the lack of convexity of the observable position (Zitterbewegung). A natural strategy to recover such positivity is to employ the well-known supersymmetric structure of the free Dirac operator, $H_{m}(0,0)=-i\boldsymbol{\alpha}\cdot\nabla+m\beta$. This means that squaring out the operator gives
$H_{m}(0,0)^{2}=(-\Delta+m^{2})\mathbb{I}_{\mathbb{C}^{4}}.$
@@
An important consequence of this spectral relation is that proving absence of eigenvalues of $H_{m}(0,0)$ reduces to prove absence of eigenvalues of second order operators of Schrödinger type for which the aforementioned positivity is guaranteed. It is easy to see that the supersymmetric property is maintained for a purely magnetic Dirac operator, namely for $H_{m}(A,0)$:
$H_{m}(A,0)^{2}=\begin{pmatrix}H_{P}+m^{2}\mathbb{I}_{\mathbb{C}^{2}}&0\\ 0&H_{P}+m^{2}\mathbb{I}_{\mathbb{C}^{2}}\end{pmatrix}$,
where $H_{P}:=-\Delta\mathbb{I}_{\mathbb{C}^{2}}+\boldsymbol{\sigma}\cdot B$ denotes the Pauli operator in $L^{2}(\mathbb{R}^{3})^{2}.$ Here $\sigma=(\sigma_{1},\sigma_{2},\sigma_{3})$ are the standard $2\times 2$ Hermitian Pauli matrices (see for instance [Cossetti2024] and [Hundertmark2024]). However, introducing a general electric potential $V$ breaks this property, making the analysis of such operators much harder. As a consequence, related results have so far been obtained only through alternative techniques; see, for instance, [DAncona2022, DAncona2022a, Dolbeault2024, Fanelli2019, Hansmann2022, Mizutani2022]. Overcoming these limitations is precisely the aim of the present work.
@@
We shall now present the main results of this paper. The first one shows that, if the potential satisfies a suitable Hardy-type smallness condition, no eigenvalues are created in the spectral gap.
@@
holds true for every $\psi\in H^{1}(\mathbb{R}^{d})$. Then $\sigma_{p}(H_{m}(A,V))=\varnothing$.
Remark 1.2.
@@
Let $d\geq 3.$ Suppose that $V:\mathbb{R}^{d}\to\mathbb{C}^{N\times N}$ is such that $V\in C^{\infty}(\mathbb{R}^{d}\setminus\{x_{0}\})$ and that $V$ is bounded outside any ball containing $x_{0}$. If $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4}\geq 0$ satisfy (1.4) and $B$ and $V$ satisfy (1.5)-(1.8), then $\sigma_{p}(H_{m}(A,V))=\varnothing.$
@@
Let $d\geq 3.$ Suppose that $V$ is as in Theorem 1.2. If $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\varepsilon_{4}$ satisfy (1.4) and $B$ and $V$ satisfy (1.5)-(1.8), then for any unitary matrix $P\in\mathbb{C}^{N\times N}$ one has $\sigma_{p}(H_{m}(A,PVP^{\ast}))=\varnothing$.
@@
This provides a sufficient condition for assumptions (1.5), (1.6) and (1.8) to be satisfied, i.e. $|W|\approx\mu|x|^{-2},$ for some suitably small coupling constant $\mu>0.$ Notice that in the specific case of our conditions, namely (1.5), (1.6) and (1.8), one has $|W|=|x|^{2}|B|^{2},$ $|W|=|x|^{2}|\nabla V|^{2},$ and $|W|=|x||\nabla V|,$ respectively.
On the other hand, one can prove (see [Ghoussoub2011, Thm. 2.1] and [Sugie2002]) that if $W$ satisfies a Hardy-type inequality (1.10), then there exist $c,R>0$ such that
@@
Finally, observe that, as a consequence of the diamagnetic inequality (see, for instance, [Lieb2001])
@@
here $\langle x\rangle:=(1+|x|^{2})^{1/2}$ is the standard notation for Japanese brackets. The alternative condition (1.11) shows that in the massive case, or if $\{\beta,V\}\neq 0$, we can indeed allow for Coulomb-type decay near the origin, but not at infinity.
@@
is included, where the potentials $\mathbb{V}_{\text{el}},\mathbb{V}_{\text{sc}},\mathbb{V}_{\text{am}}$ are customarily called electrostatic, (Lorentz-)scalar and anomalous magnetic, respectively and which are known to leave invariant the partial wave subspaces.
We emphasize that the impossibility to cover Coulomb-type potentials (1.12) in the massive case is not just technical. Indeed, it is known that, as soon as $m\neq 0,$ there are Coulomb-type perturbations that do produce discrete eigenvalues in the spectral gap (m,m)(-m,m) (see [Dolbeault2000] and references therein for the case $A=0$ and $\mathbb{V}_{\text{el}}$ and [Soff1973] for the case $\mathbb{V}_{\text{sc}}$). The situation is different for anomalous magnetic potentials $\mathbb{V}_{\text{am}}.$ In this case, since $\{\beta,\mathbb{V}_{\text{am}}\}=0,$ assumption (1.8) is no longer needed. Thus, Theorem 1.2 is valid for $V$ as in (1.12) with $\mathbb{V}=\mathbb{V}_{\text{am}}$ also in the massive case, namely for any $m\geq 0.$ We emphasize that this does not contradict the explicit eigenvalues and eigenfunctions obtained in [Cassano2018] for Dirac operators with purely anomalous magnetic perturbations. Indeed, as can be seen from the proof of their result, if the coupling constant $\delta$ (there called $\lambda$) is sufficiently small, then the corresponding eigenfunctions found there are not $L^{2}$-integrable.
@@
This is a generalization of the special case (1.12) with (1.13), and has been extensively studied as it is an interesting model in quantum mechanics (see, for example [Cassano2018] for further information).
@@
and assume (1.5), (1.6) and, if $m\neq 0,$ assume also (1.8). Moreover, assume that
@@
In the previous results we have provided a source of sufficient conditions for guaranteeing total absence of eigenvalues for $H_{A,V}.$ Nevertheless, the method of multipliers can be fruitfully used also to exclude embedded eigenvalues only, as it is shown in the following result.
Theorem 1.8.
Let $d\geq 3$. Suppose that $V:\mathbb{R}^{d}\to\mathbb{C}^{N\times N}$ is such that $V\in C^{\infty}(\mathbb{R}^{d}\setminus\{0\})$ and that $V$ is bounded outside any ball containing the origin. Assume also that there exist constants $\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\geq 0$ satisfying
$\frac{(4d-5)}{2(d-2)}\left(\frac{\varepsilon_{1}}{\varepsilon_{0}}+\varepsilon_{0}\right)+\left(\frac{4}{d-2}\right)\varepsilon_{2}\,\varepsilon_{3}+2m\varepsilon_{4}^{2}+\left(\frac{8d-6}{d-2}\right)\varepsilon_{3}+\left(\frac{4d-2}{d-2}\right)\varepsilon_{2}+\frac{4}{(d-2)^{2}}(m^{2}+1)\varepsilon_{3}^{2}\,\leq\,1,$ (1.23)
such that $V$ satisfies (1.6)-(1.8). Additionally, suppose that the following inequality
$\int|x|^{2}|B|^{2}|\psi|^{2}$ $\leq\varepsilon_{1}\|\nabla_{\!\text{A}}\psi\|^{2}+\delta\|\psi\|^{2}$ (1.24)
holds for all $\psi\in H^{1}(\mathbb{R}^{d})$ and for some $\delta\geq 0.$ Then, the Dirac operator $H_{A,V}$ has no eigenvalues on $(-\infty,-\Lambda)\cup(\Lambda,\infty)$, where
$\Lambda=\sqrt{\frac{(4d-5)}{2(d-2)}\frac{\delta}{\varepsilon_{0}}+m^{2}+1}.$ (1.25)
Remark 1.7 (Decay at infinity).
Compared to hypothesis (1.5) in Theorem 1.2, hypothesis (1.24) allows magnetic fields to decay slower. In fact, any $B$ such that $|B(x)|\leq\mu|x|^{-1}$ is admitted for $\mu$ small enough.
Remark 1.8 (Comparison with [Hundertmark2024]).
This result can be considered as a generalization of [Hundertmark2024, Thm.8.1], where purely magnetic Dirac operators were considered. More specifically, there the authors showed that under similar hypothesis to (1.24) for the magnetic field $B,$ the Dirac operator $H_{A,0}$ has no eigenvalues in $(-\infty,-\sqrt{\delta+m^{2}})\cup(\sqrt{\delta+m^{2}},\infty).$ Additionally, they showed that if $B=o(|x|^{-1}),$ then $H_{A,0}$ has no eigenvalues in $(-\infty,-m)\cup(m,\infty).$ In our case, in the limit situation $\delta=0,$ we obtain that $H_{A,V}$ does not have eigenvalues in the smaller set
$\left(-\infty,-\sqrt{m^{2}+1}\right)\cup\left(\sqrt{m^{2}+1},\infty\right).$ (1.26)
We emphasize that the translation by one in the threshold (1.26) is due to the addition of the electric potential $V.$ More precisely, as one can explicitly see from the proof of Theorem 1.8, this comes from the need of estimating
$\langle\psi,m\{\beta,V\}\psi\rangle$
for some suitable test function $\psi.$ Therefore, if $m\{\beta,V\}=0$ (in particular if $V=0$), then Theorem 1.8 shows absence of eigenvalues in $(-\infty,-m)\cup(m,\infty),$ this, in particular, recovers the result in [Hundertmark2024].
Remark 1.9.
Notice that it is possible to remove condition (1.8) from Theorem 1.8, which would allow to consider Coulomb-type potentials even in the massive case, under the cost of making smaller the interval where we can ensure the absence of eigenvalues. This is because in Theorem 1.2 condition (1.8) is used to estimate the term
$\langle 2x\cdot\nabla_{A}\psi+d\psi,m\{\beta,V\}\psi\rangle.$
And, while in Theorem 1.2 we can not allow terms containing the $L^{2}$-norm of $\psi$, we can allow them in Theorem 1.8 at a cost.
Remark 1.10.
Finally, observe that the constant $\varepsilon_{0}$ in Theorem 1.8 is not related to the smallness of the potentials. To convince oneself, it is enough to rewrite inequality (1.23) as
$\frac{(4d-5)}{2(d-2)}\frac{\varepsilon_{1}}{\varepsilon_{0}}+\left(\frac{4}{d-2}\right)\varepsilon_{2}\,\varepsilon_{3}+2m\varepsilon_{4}^{2}+\left(\frac{8d-6}{d-2}\right)\varepsilon_{3}+\left(\frac{4d-2}{d-2}\right)\varepsilon_{2}+\frac{4}{(d-2)^{2}}(m^{2}+1)\varepsilon_{3}^{2}\,\leq\,1-\frac{(4d-5)}{2(d-2)}\varepsilon_{0},$
and to note that the only purpose of the constant $\varepsilon_{0}$ is to ensure that the right hand side is positive.
@@
    We refer to [Cazacu2016] for a complete survey on the concept of magnetic field in any dimensions and its definition in terms of differential forms and tensor fields.
@@ 2. Definition of Dirac operator
In this section we gather some basic properties on the Dirac operators. We refer to [Thaller1992] for a more exhaustive dissertation.
@@
for $j,k\in\{1,2,\dots,d\}$. For an explicit construction of these matrices in any dimension see for example [Cossetti2020]. It is well known that this operator is self-adjoint in $L^{2}(\mathbb{R}^{d})$ with domain $D(H_{m})=H^{1}(\mathbb{R}^{d})$ and has pure continuous spectrum, namely
@@
for $a>0$ small enough (see [Thaller1992, Sec.4.3]). For the electric potential we shall assume that $V(x)$ is Hermitian for almost all $x\in\mathbb{R}^{d}$. Furthermore, to ensure the self-adjointness of $H_{m}(A,V)$ and, in consequence, that its spectrum is real, we will assume that $V$ is $H_{m}(A)$-bounded with relative bound less than 1. This means that there exist constants $a,b>0$ with $a\leq 1$ such that for every $\psi\in H^{1}(\mathbb{R}^{d})$ the following inequality
@@
holds. This allows us to define $H_{m}(A,V)$ as a self-adjoint operator in $L^{2}(\mathbb{R}^{d})$ with domain $H^{1}(\mathbb{R}^{d})$ by the Kato-Rellich theorem [Reed1975, Sec.X.2]. Notice that in some specific cases of special interest, condition (2.2) can be replaced by weaker assumptions.